OpenStudy (anonymous):

Solve the proportion. p/6=24/36 a. 2 c. 4 b. 3 d. 5

OpenStudy (anonymous):

@joeysampson

OpenStudy (anonymous):

@lols4life

OpenStudy (anonymous):

@mathslover

OpenStudy (jh3power):

p/6=24/36 p=6*24/36 p=6*2/3 p=4

OpenStudy (anonymous):

thanks